OpenStudy (greengalaxy):

So, I have a question from a test I took and got a 80% on it, apparently I missed this question For which operations is the set {0, 1} closed? A. Multiplication B. Division C. Addition D. Subtraction I picked A but the correct answers were A and B Why division? You can't divided 0 by 0? right?

OpenStudy (greengalaxy):

@whpalmer4 @paki @Lexaaa_ganja

OpenStudy (anonymous):

well technically 0 divided by 0 is 0...

OpenStudy (greengalaxy):

But, in the lesson they specifically said you can't do that.

OpenStudy (greengalaxy):

@satellite73 @Gabriel_Morrow @Hero

OpenStudy (greengalaxy):

@helper9000 @pooja195 Anyone?

OpenStudy (whpalmer4):

@mrqueest 0 / 0 is NOT 0 For all values of \(a,b,c\), if \[a/b = c\] then \[c*b =a\] Does that work for \(a/0=0\)? \[3/0 = 0\]means that \[0*0 = 3\]but it isn't.

OpenStudy (greengalaxy):

Oh! Wow, thank you! I wouldn't have gotten this at all! Thanks!

OpenStudy (whpalmer4):

My guess is that division was included with the understanding that division by 0 is excluded because it gives an undefined result.

OpenStudy (greengalaxy):

I gave you a medal, also im closing this session, thanks a ton!

OpenStudy (anonymous):

yes green?

OpenStudy (greengalaxy):

It's alright my question is already awnserd.Thanks anyways

OpenStudy (anonymous):

ok np @GreenGalaxy